OpenStudy (anonymous):

What would the ordered pair be for the following system of equations: x + 2y = 4 (1) x = 7-2y (2)

OpenStudy (anonymous):

x+2y=4 x=4-2y x=7-2y 7-2y=4-2y since 4 is not equal to 7 it has no solution.

OpenStudy (anonymous):

1 => x+2y = 4 2 => x+2y = 7 which cannot both be true at the same time